Obtain the 6th root of (-7)
1
Expert's answer
2020-06-08T16:56:36-0400

"-7=7(-1+0*i)=7e^{i\pi}"

"\sqrt[6]{7e^{i\pi}}=\sqrt[6]{7}e^{i\frac{\pi+2\pi k}{6}},k=0..5"